# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 3.3 Alignment of maxima, equivalent ripples
An obvious consequence of Eq. (87) is the configuration where the leading order off-set is absent. Namely, coupled set (87) reduces to two coupled equations in case all trapping domain are equivalent which means that all relevant maxima are aligned with a global maximum, $x_{j}^{\pm}=0$, and all aspect ratios are the same (equal to one), $A_{j}=1$. First condition means that global maximum is reached on a line (lines) rather than at a point (points). Second condition means that all $\delta\eta_{j}$ are the same, i.e. integral (78) between maximum points is the same for all trapping domains. In other words, this means that the value of the first integral (26) at the trapped-passing boundary $\eta=\eta_{b}$ does not depend on field line parameter $\vartheta_{0}$,
@@
Obviously, deep in the trapped region $x\rightarrow-\infty$ where $\Phi(x)\rightarrow 0$ and $\alpha^{+}\rightarrow const$, this equation reduces to $\alpha^{+}=-\alpha^{+}$, i.e. it results in zero off-set $\alpha^{\pm}=0$. Actually, as shown in (Helander et al., 2011), ripple equivalence condition (89) which is fulfilled for a perfectly quasi-isodynamic stellarator where the magnetic field is omnigeneous, allows to reduce the boundary layer problem to a tokamak like problem up to the same degree of accuracy as the analysis here (essentially, we use the integral representation of the same problem as in (Helander et al., 2011)).
@@
Figure 11: Normalized distribution functions $\alpha^{+}$ (blue), $\alpha^{-}$ (red) and source function $\Phi$ (magenta). Full normalized distribution functions, $\alpha^{\pm}+\alpha_{1/\nu}^{\pm}$, are shown with dashed lines. Right picture is a zoom to the trapped region, where also the asymptotic solution (92) is shown (black dashed).
@@
where $\alpha_{0}=\text{const}$. They agree well because the latter solution must actually be the same with $\alpha^{\pm}$ in the region where the source term is negligible. We can recover the full solution if we add to $\alpha^{\pm}(x)$ the rescaled leading order term (45) expanded to the linear order in $\eta-\eta_{b}$,
@@
Figure 12: Normalized distribution functions $\alpha_{j}^{+}$ (solid) and $\alpha_{j}^{-}$ (dashed) in case of three relevant trapping domains per closed field line. Right picture is a zoom to the trapped particle domain.
A finite off-set in the first and last domains, $\alpha_{2}^{\pm}(-\infty)=-\alpha_{0}^{\pm}(-\infty)$ can be clearly seen. In such a case, function $g_{\text{bou}}$ in the trapped particle domain and, respectively, bootstrap coefficient diverge at low collisionality as $\lambda_{bB}\sim\nu^{-1/2}$ (see the normalization (76) and (86)).
@@
Figure 13: Normalized bootstrap coefficient $\lambda_{bB}$, Eq. (43), for “anti-sigma optimized” configurations as function of normalized collisionality $\nu_{\ast}$. The right picture is a log-log plot with $\nu^{-1/2}$ trend added (black dotted). Cases 1 and 2 correspond to non-equivalent ($\varepsilon_{3}=0$) and equivalent ripples, respectively, and $\iota=1/4$. Case 3 corresponds to non-equivalent ripples and $\iota=(1+\sqrt{5})/20$. Shaing-Callen limit (42) is shown with dashed lines.
@@
with $\varepsilon_{0}=0.125$, $\varepsilon_{1}=0.05$ and two values of $\varepsilon_{3}$. We set $N_{\rm tor}=1$ and $\iota=1/4$ to get 4 relevant domains (ripples) on a closed field line. In case $\varepsilon_{3}=0$, condition $A_{j}=1$ does not hold, and a strong off-set of bootstrap coefficient $\lambda_{bB}\propto\nu^{-1/2}$ is seen in Fig. 13. In case $\varepsilon_{3}=0.06994425$ all four ripples are equivalent, $A_{j}=1$, and the leading order off-set is absent. However, there still remains a residual off-set driven by source $\Delta s$, Eq. (71), which includes a constant and logarithmic parts, $\delta\lambda_{bB}\rightarrow C_{1}+C_{2}\log(\nu_{\ast})$, see the previous section. It should be noted that $\nu^{-1/2}$ divergence of $\lambda_{bB}$ in case of aligned maxima but non-equivalent ripples is not a property of a closed field line. A similar behavior is seen also for the case $\varepsilon_{3}=0$ at the irrational flux surface with $\iota=(1+\sqrt{5})/20\approx 0.1618$ labelled there as case 3.
@@ 3.4 Non-aligned maxima, large aspect ratio
Let us consider now a more typical situation where local field maxima are not aligned. For simplicity, we consider a configuration with short ripples located near global maximum adjacent to a long ripple, as in the case of the first rippled tokamak configuration (59) with $\iota=1/4$ illustrated in Figs. 3 and 6. Besides the global maximum, there are only two relevant maxima at low collisionalities. They separate two short and shallow trapping domains called below “off-set domains” from a wide domain in between called below “main region”. Setting $\delta\eta_{\rm ref}=\delta\eta_{1}$ so that $A_{1}=1$, aspect ratios for off-set domains are the same, $A_{0}=A_{2}\equiv A_{o}$. Respective value $A_{o}\approx 23.7\gg 1$ will be called below “off-set aspect ratio” or simply “aspect ratio”. To keep an example more general than configuration with $\iota=1/4$, we do not assume that local maxima separating the main region from two off-set domains are the same (thus allowing broken stellarator symmetry). For the present example, it is convenient to change general notation in Eqs. (87) as follows, $\alpha_{0}^{\pm}\equiv\alpha_{-o}^{\pm}$, $\alpha_{1}^{\pm}\equiv\alpha^{\pm}_{mr}$ and $\alpha_{2}^{\pm}\equiv\alpha_{+o}^{\pm}$ so that subscripts “$+o$” and “$-o$” denote the left and right off-set domain, respectively, and subscript “$mr$” stands for the main region. Respective normalized mis-alignments of local maxima we re-notate as follows, $x_{1}^{+}=x_{2}^{-}\equiv x_{-o}$ and $x_{3}^{-}=x_{2}^{+}\equiv x_{+o}$. Thus, we cast the closed set of six equations (87) to
@@
Obviously, in the leading order over $A_{o}^{-1}$, these equations are the same with Eqs. (90) for equivalent ripples.
@@
Substituting this result in Eq. (97) we note also that variation of the last term in Eq. (97) is determined in the region of interest by $\tilde{\Theta}$ while the factor $\alpha_{ma}^{\pm}$ varies slowly and can be replaced by its value at $x=x_{\pm o}$. Introducing the new independent variable $y=A_{o}(x-x_{\pm o})$, equation set for $\alpha^{\pm}_{\pm o}$ reduces to two independent equations
@@
The last conclusion can be easily generalized to the case where a short off-set domain is separated from two long main regions by local maxima on both sides. Assuming that solutions in main regions are known, the off-set is determined by the value of distribution function in the main region separated from the off-set domain by lower local maximum which determines the class transition boundary limiting the phase space region for particles locally trapped in the off-set well.
@@
Since $g_{\text{bou}}$ is given there by propagator method in the leading order only, we added error bars to the result of Eq. (87) showing the order of magnitude estimate of the omitted correction term $\Delta g_{\text{bou}}$ driven by the source $\Delta s$ in Eq. (69). Roughly estimating $\Delta g_{\text{bou}}\sim g_{0}^{t}$ at the trapped passing boundary (function $\Delta g_{\text{bou}}$ must balance $g_{0}^{t}$ for restoring the periodicity in the passing region), we estimate it further in the trapped particle domain as $\Delta g_{\text{bou}}\sim g_{0}^{t}\alpha^{+}(x)$ where $\alpha^{+}(x)$ is solution for equivalent ripples (Eq. (91)) and $x$ is defined in Eq. (86).
@@ 3.5 Weakly non-equivalent ripples
Corresponding dimensional off-sets via (86) can be combined to
@@ 3.6 Off-set of Ware pinch coefficient
We define the off-set in the normalized Ware pinch (bootstrap) coefficient as follows, $\lambda_{\text{off}}\equiv\lambda_{bB}-\lambda_{bB}^{\infty}$, where $\lambda_{bB}=\lambda_{bB}^{\dagger}$ is defined by Eq. (57) via mono-energetic $\bar{D}_{13}$ for finite collisionality and $\lambda_{bB}^{\infty}$ denotes its asymptotic ($l_{c}\rightarrow\infty$) value (42). Thus, $\lambda_{\text{off}}$ is given by Eq. (57) with the replacement $g_{(3)}\rightarrow g_{\text{off}}=g_{(3)}-g_{(3)}^{\infty}$ in the definition (11) of $\bar{D}_{13}$, where $g_{(3)}^{\infty}$ is the asymptotic solution $g_{0}$ derived in Section 2.2. Splitting in (11) the integration interval $[\varphi_{0},\varphi_{N}]$ into sub-intervals $[\varphi_{j},\varphi_{j+1}]$ separated by relevant maxima $\varphi_{j}$ (points of highest local maxima contained within the matching boundary, $\eta_{m}B(\varphi_{j})>0$, see Section 3.2), these intervals correspond to off-set domains or main regions where the even part of the distribution function $g_{\text{off}}$ is essentially different from zero in the trapped particle domain and close to a constant $g_{\text{off}}^{(j)}$ in each interval $\varphi_{j}<\varphi<\varphi_{j+1}$, i.e. $g_{\text{off}}(\varphi,\eta)\approx\Theta(\eta-\eta_{b})g_{\text{off}}^{(j)}$, we obtain the off-set coefficient as
$\lambda_{\text{off}}\approx\sum_{j}g_{\text{off}}^{(j)}w_{\text{off}}^{(j)},$ (109)
where weighting factors are
$w_{\text{off}}^{(j)}=\frac{3}{2\rho_{L}B}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}\int\limits_{\varphi_{j}}^{\varphi_{j+1}}{\rm d}\varphi\int\limits_{\eta_{\text{loc}}^{(j)}}^{1/B}{\rm d}\eta\;s_{(1)}^{\dagger}=-\frac{3}{2\rho_{L}B}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}H_{j}(\eta_{\text{loc}}),$ (110)
$H_{j}(\eta)$ is given by (26), $\eta_{\text{loc}}^{(j)}=\max(\eta_{j},\eta_{j+1})$, and we assume here the closed field line (i.e. finite $\varphi_{N}$).
@@
$w_{\text{off}}^{(j)}=-\frac{1}{2}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}\int\limits_{\varphi_{j}}^{\varphi_{j+1}}{\rm d}\varphi\frac{|\nabla r|k_{G}}{BB^{\varphi}}\left(4-\eta_{\text{loc}}^{(j)}B\right)(1-\eta_{\text{loc}}^{(j)}B)^{1/2}.$ (113)
@@
$w_{\text{off}}^{(j)}=\frac{{\rm d}r}{{\rm d}\psi}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{2}}\right)^{-1}I_{{\cal V}}^{(j)},\qquad I_{{\cal V}}^{(j)}=\int\limits_{\varphi_{j}}^{\varphi_{j+1}}{\rm d}\varphi{\cal V},$ (114)
@@
For configuration with $\iota=1/4$, off-set is fully determined by the first (and last) off-set well, $g^{(0)}_{\text{off}}$, since $g^{(1)}_{\text{off}}=0$ and $g^{(2)}_{\text{off}}=-g^{(0)}_{\text{off}}$ due to stellarator symmetry. For configuration with $\iota=2/5$, both, the off-set well, $g^{(0)}_{\text{off}}$, and the main region, $g^{(1)}_{\text{off}}$, are contributing (with $g^{(3)}_{\text{off}}=-g^{(1)}_{\text{off}}$ and $g^{(4)}_{\text{off}}=-g^{(0)}_{\text{off}}$ just doubling the result due to stellarator symmetry). Off-set wells merge with main regions (0 with 1 and 3 with 4) at $\nu_{\ast}<10^{-4}$, therefore we used $\eta_{\text{loc}}^{(j)}$ for such combined regions in individual $w^{(j)}_{\text{off}}$. It can be seen that the contribution of the off-set wells is slightly larger than of the main regions, despite much smaller toroidal extent. These two contributions tend to compensate each other, which is natural for bounce-averaged drift (the longer the orbit, the smaller is this drift).

### v3

@@ 3.3 Alignment of maxima, equivalent ripples
An obvious consequence of Eq. (83) is the configuration where the leading order off-set is absent. Namely, coupled set (83) reduces to two coupled equations in case all trapping domain are equivalent which means that all relevant maxima are aligned with a global maximum, $x_{j}^{\pm}=0$, and all aspect ratios are the same (equal to one), $A_{j}=1$. The first condition means that global maximum is reached on a line (lines) rather than at a point (points). The second condition means that all $\delta\eta_{j}$ are the same, i.e. the integral (74) between maximum points is the same for all trapping domains. In other words, this means that the value of the first integral (26) at the trapped-passing boundary $\eta=\eta_{b}$ does not depend on field line parameter $\vartheta_{0}$,
@@
Obviously, deep in the trapped region $x\rightarrow-\infty$ where $\Phi(x)\rightarrow 0$ and $\alpha^{+}\rightarrow const$, this equation reduces to $\alpha^{+}=-\alpha^{+}$, i.e. it results in zero off-set $\alpha^{\pm}=0$. Actually, as shown in (Helander et al., 2011), the ripple equivalence condition (85) which is fulfilled for a perfectly quasi-isodynamic stellarator where the magnetic field is omnigeneous, allows reducing the boundary layer problem to a tokamak like problem up to the same degree of accuracy as the analysis here (essentially, we use the integral representation of the same problem as in (Helander et al., 2011)).
@@
Figure 12: Normalized distribution functions $\alpha^{+}$ (blue), $\alpha^{-}$ (red) and source function $\Phi$ (magenta). Full normalized distribution functions, $\alpha^{\pm}+\alpha_{1/\nu}^{\pm}$, are shown with dashed lines. The right picture is a zoom to the trapped region, where also the asymptotic solution (88) is shown (black dashed).
@@
where $\alpha_{0}=\text{const}$. They agree well because the latter solution must actually be the same as $\alpha^{\pm}$ in the region where the source term is negligible. We can recover the full solution if we add to $\alpha^{\pm}(x)$ the rescaled leading order term (45) expanded to the linear order in $\eta-\eta_{b}$,
@@
Figure 13: Normalized distribution functions $\alpha_{j}^{+}$ (solid) and $\alpha_{j}^{-}$ (dashed) in case of three relevant trapping domains per closed field line. The right picture is a zoom to the trapped particle domain.
A finite off-set in the first and last domains, $\alpha_{2}^{\pm}(-\infty)=-\alpha_{0}^{\pm}(-\infty)$ can be clearly seen. In such a case, the function $g_{\text{bou}}$ in the trapped particle domain and, respectively, bootstrap coefficient diverge at low collisionality as $\lambda_{bB}\sim\nu^{-1/2}$ (see the normalization (72) and (82)).
@@
Figure 14: Left: Normalized bootstrap coefficient $\lambda_{bB}$, Eq. (43), computed by NEO-2 for “anti-sigma optimized” configurations in cases 1, 2 and 3 (see the legend) as function of normalized collisionality $\nu_{\ast}$. For each case, the Shaing-Callen limit (42) is shown with a dashed line of the respective color. Sign of $\lambda_{bB}$ (but not of the respective asymptotic) is reversed in case 2. Right: Magnetic field strength $B(\varphi)$, Eq. (90) for these cases plotted along the field line within the first four toroidal periods.
@@
with $\varepsilon_{0}=0.125$, $\varepsilon_{1}=0.05$ and two values of $\varepsilon_{3}$. We set $N_{\rm tor}=1$ and $\iota=1/4$ to get 4 relevant domains (ripples) on a closed field line passing through the point $(\vartheta,\varphi)=(0,0)$. In case $\varepsilon_{3}=0$ (case 1), condition $A_{j}=1$ does not hold, and a strong off-set of bootstrap coefficient $\lambda_{bB}\propto\nu^{-1/2}$ is seen in Fig. 14. In case $\varepsilon_{3}=0.06994425$ (case 2) all four ripples are equivalent, $A_{j}=1$, and the leading order off-set is absent. However, there still remains a residual off-set driven by source $\Delta s$, Eq. (67), which includes a constant and logarithmic parts, $\delta\lambda_{bB}\rightarrow C_{1}+C_{2}\log(\nu_{\ast})$, see the previous section. It should be noted that $\nu^{-1/2}$ divergence of $\lambda_{bB}$ in case of aligned maxima but non-equivalent ripples is not a property of a closed field line. A similar behavior is seen also for the case $\varepsilon_{3}=0$ at the irrational flux surface with $\iota=(1+\sqrt{5})/20\approx 0.1618$ labelled there as case 3.
@@ 3.4 Non-aligned maxima, large aspect ratio
Let us consider now a more typical situation where local field maxima are not aligned. For simplicity, we consider a configuration with short ripples located near global maximum adjacent to a long ripple, as in the case of the first rippled tokamak configuration (55) with $\iota=1/4$ illustrated in Figs. 3 and 6. Besides the global maximum, there are only two relevant maxima at low collisionalities. They separate two short and shallow trapping domains called below “off-set domains” from a wide domain in between called below “main region”. Setting $\delta\eta_{\rm ref}=\delta\eta_{1}$ so that $A_{1}=1$, aspect ratios for off-set domains are the same, $A_{0}=A_{2}\equiv A_{o}$. Respective value $A_{o}\approx 23.7\gg 1$ will be called below “off-set aspect ratio” or simply “aspect ratio”. To keep an example more general than configuration with $\iota=1/4$, we do not assume that local maxima separating the main region from two off-set domains are the same (thus allowing broken stellarator symmetry). For the present example, it is convenient to change general notation in Eqs. (83) as follows, $\alpha_{0}^{\pm}\equiv\alpha_{-o}^{\pm}$, $\alpha_{1}^{\pm}\equiv\alpha^{\pm}_{mr}$ and $\alpha_{2}^{\pm}\equiv\alpha_{+o}^{\pm}$ so that subscripts “$+o$” and “$-o$” denote the left and right off-set domain, respectively, and subscript “$mr$” stands for the main region. Respective normalized mis-alignments of local maxima we re-notate as follows, $x_{1}^{+}=x_{2}^{-}\equiv x_{-o}$ and $x_{3}^{-}=x_{2}^{+}\equiv x_{+o}$, see schematic Fig. 15. Thus, we cast the closed set of six equations (83) to
@@
Figure 15: Locations of normalized outgoing particle distributions $\alpha_{\pm o}^{\pm}$ and $\alpha_{mr}^{\pm}$ in the respective off-set domains and the main region. Positions of relevant maxima $\varphi_{j}$ are shown with blue lines. Normalized misalignments $x_{\pm o}$ are shown with red lines.
@@
Obviously, in the leading order over $A_{o}^{-1}$, these equations are the same as Eqs. (86) for equivalent ripples.
@@
Substituting this result in Eq. (93) we note also that variation of the last term in Eq. (93) is determined in the region of interest by $\tilde{\Theta}$ while the factor $\alpha_{mr}^{\pm}$ varies slowly and can be replaced by its value at $x=x_{\pm o}$. Introducing the new independent variable $y=A_{o}(x-x_{\pm o})$, equation set for $\alpha^{\pm}_{\pm o}$ reduces to two independent equations
@@
The last conclusion can be easily generalized to the case where a short off-set domain is separated from two long main regions by local maxima on both sides. Assuming that solutions in both main regions are known, the off-set is determined by the solution in the main region separated from the off-set domain by the lower local maximum (this local maximum determines the class transition boundary $\eta_{c}$ limiting the region $\eta>\eta_{c}$ with particles trapped in the off-set domain).
@@
Since $g_{\text{bou}}$ is given there by propagator method in the leading order only, we added error bars to the result of Eq. (83) showing the order of magnitude estimate of the omitted correction term $\Delta g_{\text{bou}}$ driven by the source $\Delta s$ in Eq. (65). According to Eq. (66), this correction term is driven by the sources in the passing particle region and is of the order of $g_{0}^{t}$ at the trapped passing boundary (and in the whole boundary layer) because $\Delta g_{\text{bou}}$ must restore periodicity in the passing region. Since there are no sources in the trapped particle region, $\Delta g_{\text{bou}}$ decays exponentially outside the boundary layer, as follows from Eq. (62) and from the solution $\alpha^{+}(x)$ for equivalent ripples, Eq. (87), where $x$ is defined in Eq. (82). Therefore, we extend the estimate of the correction term to the whole trapped particle region by using $\Delta g_{\text{bou}}\sim g_{0}^{t}\alpha^{+}(x)$. This simple estimation gives an error bar $g_{0}^{t}\alpha^{+}(x_{\pm o})$ that decreases when $x_{\pm o}\propto\nu_{\ast}^{-1/2}$ increases.
@@ 3.5 Weakly non-equivalent ripples
The corresponding dimensional off-set for the ripple $j$ related in the general case to the dimensionless off-set via (82) as follows,
$g_{\text{off}}^{(j)}\equiv g_{j(out)}^{+}(\infty)=g_{j(out)}^{-}(\infty)=\frac{C_{0}\delta\eta_{\text{ref}}}{\sqrt{8}}\alpha_{j}^{\pm}(-\infty),$ (104)
is expressed in the present case via the off-set factors as
@@ 3.6 Off-set of Ware pinch coefficient
We define the off-set in the normalized Ware pinch (bootstrap) coefficient as follows, $\lambda_{\text{off}}\equiv\lambda_{bB}-\lambda_{bB}^{\infty}$, where $\lambda_{bB}=\lambda_{bB}^{\dagger}$ is defined by Eq. (53) via mono-energetic $\bar{D}_{13}$ for finite collisionality and $\lambda_{bB}^{\infty}$ denotes its asymptotic ($l_{c}\rightarrow\infty$) value (42). Thus, $\lambda_{\text{off}}$ is given by Eq. (53) with the replacement $g_{(3)}\rightarrow g_{\text{off}}=g_{(3)}-g_{(3)}^{\infty}$ in the definition (11) of $\bar{D}_{13}$, where $g_{(3)}^{\infty}$ is the asymptotic solution $g_{0}$ derived in Section 2.2. Splitting in (11) the integration interval $[\varphi_{0},\varphi_{N}]$ into sub-intervals $[\varphi_{j},\varphi_{j+1}]$ separated by relevant maxima $\varphi_{j}$ (points of highest local maxima $\eta_{j}B(\varphi_{j})=1$ contained within the matching boundary $\eta_{m}$, $\eta_{j}<\eta_{m}$, see Section 3.2), these intervals correspond to off-set domains or main regions where the even part of the distribution function $g_{\text{off}}$ is essentially different from zero in the trapped particle domain and is close to a constant $g_{\text{off}}^{(j)}$ approximated using propagator techniques by Eq. (104) in each interval $\varphi_{j}<\varphi<\varphi_{j+1}$ where we set $g_{\text{off}}(\varphi,\eta)\approx\Theta(\eta-\eta_{\text{loc}}^{(j)})\Theta(\eta B(\varphi)-1)\;g_{\text{off}}^{(j)}$ with $\eta_{\text{loc}}^{(j)}=\max(\eta_{j},\eta_{j+1})$, we obtain the off-set coefficient as
$\lambda_{\text{off}}\approx\sum_{j}g_{\text{off}}^{(j)}w_{\text{off}}^{(j)}.$ (106)
Here, weighting factors are
$w_{\text{off}}^{(j)}=\frac{3}{2\rho_{L}B}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}\int\limits_{\varphi_{(j)}^{-}}^{\varphi_{(j)}^{+}}{\rm d}\varphi\int\limits_{\eta_{\text{loc}}^{(j)}}^{1/B}{\rm d}\eta\;s_{(1)}^{\dagger}=-\frac{3}{2\rho_{L}B}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}H_{j}\left(\eta_{\text{loc}}^{(j)}\right),$ (107)
$H_{j}(\eta)$ is given by (26), $\varphi_{(j)}^{\pm}=\varphi_{j}^{\pm}(\eta_{\text{loc}}^{(j)})$ are coordinates of banana tips (turning points, see the definition of $\varphi_{j}^{\pm}(\eta)$ below Eq. (14)) of the transient orbit separating locally trapped orbits in the offset well from the rest phase space (one of tips $\varphi_{(j)}^{\pm}$ is located at the point of lowest local maximum, $\varphi_{j}$ or $\varphi_{j+1}$), and we assume here the closed field line (i.e. finite $\varphi_{N}$).
@@
$w_{\text{off}}^{(j)}=-\frac{1}{2}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{\varphi}}\right)^{-1}\int\limits_{\varphi_{(j)}^{-}}^{\varphi_{(j)}^{+}}{\rm d}\varphi\frac{|\nabla r|k_{G}}{BB^{\varphi}}\left(4-\eta_{\text{loc}}^{(j)}B\right)(1-\eta_{\text{loc}}^{(j)}B)^{1/2}.$ (110)
@@
$w_{\text{off}}^{(j)}=\frac{{\rm d}r}{{\rm d}\psi}\left(\int\limits_{\varphi_{0}}^{\varphi_{N}}\frac{{\rm d}\varphi}{B^{2}}\right)^{-1}I_{{\cal V}}^{(j)},\qquad I_{{\cal V}}^{(j)}=\int\limits_{\varphi_{(j)}^{-}}^{\varphi_{(j)}^{+}}{\rm d}\varphi{\cal V},$ (111)
@@
For configuration with $\iota=1/4$, off-set is fully determined by the first (and last) off-set well, $g^{(0)}_{\text{off}}$, since $g^{(1)}_{\text{off}}=0$ and $g^{(2)}_{\text{off}}=-g^{(0)}_{\text{off}}$ due to stellarator symmetry. For configuration with $\iota=2/5$, both, the off-set well, $g^{(0)}_{\text{off}}$, and the main region, $g^{(1)}_{\text{off}}$, are contributing (with $g^{(3)}_{\text{off}}=-g^{(1)}_{\text{off}}$ and $g^{(4)}_{\text{off}}=-g^{(0)}_{\text{off}}$ just doubling the result due to stellarator symmetry). Off-set wells merge with main regions (0 with 1 and 3 with 4) at $\nu_{\ast}<10^{-4}$, therefore we used $\eta_{\text{loc}}^{(j)}$ for such combined regions in individual $w^{(j)}_{\text{off}}$. It can be seen that the contribution of the off-set wells is slightly larger than of the main regions, despite much smaller toroidal extent, which is natural for bounce-averaged drift (the longer the orbit, the smaller is this drift). These two contributions tend to compensate each other at lower collisionalities, which is a general trend due to the respective increase of the off-set domain length studied in more detail in Section 4.2.